# Mechanistic Interpretability for AI Safety A Review

### v2

@@ 4 Core Methods
Mechanistic interpretability employs tools and techniques adopted from various interpretability approaches, focusing on causal methods that distinguish it from traditional, more observational techniques. This section provides an overview of the essential methodologies, enabling detailed observation and analysis of neural network models (Section 4.1), as well as interventional methods that allow for direct manipulation within the model (Section 4.2). The interplay between observation and intervention facilitates a comprehensive understanding of neural network operations (Section 4.3). Figure 7 provides an overview of the relevant methods and techniques.
Figure 7: Overview of relevant methods and techniques employed in mechanistic interpretability research. Observational methods proposed for mechanistic interpretability include structured probes (more aligned with top-down interpretability), logit lens variants, and sparse autoencoders (SAEs). Additionally, as mechanistic interpretability focuses on causal understanding, novel methods encompass variants of activation patching for uncovering causal mechanisms and causal scrubbing for hypothesis evaluation.
@@ 4.1 Observation
  • Example-based methods identify real input examples that highly activate specific neurons or layers. This helps pinpoint influential data points that maximize neuron activation within the neural network.
  • Feature-based methods encompass techniques that generate synthetic inputs to optimize neuron activation. These neuron visualization techniques reveal how neurons respond to stimuli and which features are sensitive to (Zeiler & Fergus, 2014). By understanding the synthetic inputs that drive neuron behavior, we can hypothesize about the features encoded by those neurons.
@@
Probing involves training a classifier using the activations of a model, with the classifier’s performance subsequently observed to deduce insights about the model’s behavior and internal representations. As highlighted by Belinkov (2021), this technique faces a notable challenge: the probe’s performance may often reflect its own learning capacities more than the actual characteristics of the model’s representations. This dilemma has led researchers to investigate the ideal balance between the complexity of a probe and its capacity to accurately represent the model’s features (Cao et al., 2021; Voita & Titov, 2020).
The linear representation hypothesis offers a resolution to this issue. Under this hypothesis, the failure of a simple linear probe to detect certain features suggests their absence in the model’s representations. Conversely, if a more complex probe succeeds where a simpler one fails, it implies that the model contains features that a complex function can combine into the target feature. Still, the target feature itself is not explicitly represented. This hypothesis implies that using linear probes could suffice in most cases, circumventing the complexity considerations generally associated with probing (Belinkov, 2021).
A significant limitation of probing is the inability to draw behavioral or causal conclusions. The evidence provided by probing is mainly observational, focusing on what information is encoded rather than how it is used (also see Figure 1). This necessitates careful analysis and possibly the adoption of alternative approaches (Elazar et al., 2021) or the integration of intervention techniques to draw more substantive conclusions about the model’s behavior (Section 4.2).
Probing has been used to analyze the acquisition of chess knowledge in AlphaZero (McGrath et al., 2022) and the representation of linguistic information in BERT (Tenney et al., 2019). Gurnee et al. (2023) introduce sparse probing, decoding internal neuron activations in large models to understand feature representation and sparsity. They show that early layers use sparse combinations of neurons to represent many features in superposition, while middle layers have dedicated monosemantic neurons for higher-level contextual features.
@@
Structured probes aid conceptual interpretability, probing language models for complex features like truth representations. Notably, Burns et al. (2023)’s contrast-consistent search (CCS) method identifies linear projections exhibiting logical consistency in hidden states, contrasting truth values for statements and negations.
However, structured probes face significant challenges in unsupervised probing scenarios. As Farquhar et al. (2023) showed, arbitrary features, not just knowledge-related ones, can satisfy the CCS loss equally well, raising doubts about scalability. For example, the loss may capture simulation of knowledge from hypothesized simulacra within sufficiently powerful language models rather than the models’ true knowledge. Furthermore, Farquhar et al. (2023) demonstrates unsupervised methods like CCS often detect prominent but unintended features in the data, such as distractors like "banana." The discovered features are also highly sensitive to prompt choice, and there is no principled way to select prompts that would reliably surface a model’s true knowledge.
@@
[htb] Sparse Dictionary Learning Sparse autoencoders (Cunningham et al., 2024) represent a solution attempt to the challenge of polysemantic neurons. The problem of superposition is mathematically formalized as sparse dictionary learning (Olshausen & Field, 1997) problem to decompose neural network activations into disentangled component features. The goal is to learn a dictionary of vectors $\{\mathbf{f}_{k}\}_{k=1}^{n_{\text{feat}}}\subset\mathbb{R}^{d}$ that can represent the unknown, ground truth network features as sparse linear combinations. If successful, the learned dictionary contains monosemantic neurons corresponding to features (Sharkey et al., 2022b). The autoencoder architecture consists of an encoder and a ReLU activation function, expanding the input dimensionality to $d_{\text{hid}}>d_{\text{in}}$. The encoder’s output is given by: $\mathbf{h}$ $=\text{ReLU}(W_{\text{enc}}\mathbf{x}+\mathbf{b})$ (1) ${\mathbf{x^{\prime}}}$ $=W_{\text{dec}}\mathbf{h}=\sum_{i=0}^{d_{\text{hid}}-1}h_{i}\mathbf{f}i$ (2) where $W_{\text{enc}},W_{\text{dec}}^{T}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{in}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{hid}}}$. The parameter matrix $W_{\text{dec}}$ forms the feature dictionary, with rows $\mathbf{f}_{i}$ as dictionary features. The autoencoder is trained to minimize the loss, where the ${L}_{1}$ penalty on $\mathbf{h}$ encourages sparse reconstructions using the dictionary features. $\mathcal{L}(\mathbf{x})=||\mathbf{x}-{\mathbf{x^{\prime}}}||_{2}^{2}+\alpha||\mathbf{h}||_{1}$ (3) [Uncaptioned image] Figure 8: Illustration of a sparse autoencoder applied to the MLP layer activations, consisting of an encoder that increases dimensionality while emphasizing sparse representations and a decoder that reconstructs the original activations using the learned feature dictionary.
@@
As highlighted in Section 3.1, recent work suggests that the essential elements in neural networks are linear combinations of neurons representing features in superposition (Elhage et al., 2022b). Sparse autoencoders provide a methodology to decompose neural network activations into these individual component features (Sharkey et al., 2022b; Cunningham et al., 2024). This process involves reconstructing activation vectors as sparse linear combinations of directional vectors within the activation space, a problem also known as sparse dictionary learning (Olshausen & Field, 1997).
Sparse dictionary learning has led to the development of various sparse coding algorithms (Lee et al., 2006). The sparse autoencoder stands out for its simplicity and scalability (Sharkey et al., 2022b). The first application to a language model was by Yun et al. (2021), who implemented sparse dictionary learning across multiple layers of a language model.
Sparse autoencoders, a variant of the standard autoencoder framework, incorporate sparsity regularization to encourage learning sparse yet meaningful data representations. Theoretical foundations in the disentanglement literature suggest that autoencoders can recover ground truth features under feature sparsity and non-negativity (Whittington et al., 2022). The "ground truth features" here refer to the true, disentangled features that underlie the data distribution, which the autoencoder aims to recover through its sparse encoding. In the context of neural networks, these would correspond to the individual features combined to form neuron activations, which the sparse autoencoder attempts to disentangle and represent explicitly in its dictionary.
Practical implementations, such as the toy model by Sharkey et al. (2022b), demonstrate the viability of this approach, with the precise tuning of the sparsity penalty on the hidden activations being a critical aspect that dictates the sparsity level of the autoencoder. We show an overview in the pink box on sparse autoencoders in Figure 8.
Empirical studies indicate that sparse autoencoders can enhance the interpretability of neural networks, exhibiting higher scores on the autointerpretability metric and increased monosemanticity (Bricken et al., 2023; Cunningham et al., 2024; Sharkey et al., 2022b). Furthermore, sparse autoencoders have been employed to measure feature sparsity (Deng et al., 2023) and interpret reward models in reinforcement learning-based language models (Marks et al., 2023a), making them an actively researched area in mechanistic interpretability.
@@ 4.2 Intervention
The theory of causality (Pearl, 2009) provides a mathematically precise framework for mechanistic interpretability, offering a rigorous approach to understanding high-level semantics in neural representations (Geiger et al., 2023a). By treating neural networks as causal models, with their compute graphs serving as causal graphs, researchers can perform precise interventions and examine the roles of individual parameters (McGrath et al., 2023). This causal perspective on interpretability has led to the development of various intervention techniques, including activation patching (Section 4.2.1), causal abstraction (Section 4.2.2), and hypothesis testing methods (Section 4.2.3).
@@ 4.2.1 Activation Patching
Figure 9: (a) Activation patching in a transformer model. Left: The model processes the clean input "Colosseum in Rome," caching the latent activations (step i). Right: The model runs with the corrupted input "Eiffel Tower in Paris" (step ii). The pink arrow shows an MLP layer activation (green diamond) patched from the clean run into the corrupted run (step iii). This causes the prediction to change from "Paris" to "Rome," demonstrating how the significance of the patched component is determined (step iv). By comparing these carefully selected inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the location prediction behavior. (b) Activation patching directions: Top: Patching corrupted activations (orange) into clean circuits (turquoise) reveals sufficient components for identifying OR logic scenarios. Bottom: Patching clean activations (green) into corrupted circuits (orange) reveals necessary components that are useful for identifying AND logic scenarios. The AND and OR gates demonstrate how these patching directions uncover different logical relationships between model components.
@@
The standard protocol for activation patching (Figure 9a) involves: (step i) running the model with a clean input and caching the latent activations; (step ii) executing the model with a corrupted input; (step iii) re-running the model with the corrupted input but substituting specific activations with those from the clean cache; and (step iv) determining significance by observing the variations in the model’s output during the third step, thereby highlighting the importance of the replaced components. This process relies on comparing pairs of inputs: a clean input, which triggers the desired behavior, and a corrupted input, which is identical to the clean one except for critical differences that prevent the behavior. By carefully selecting these inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the behavior.
Differences in patching direction—clean to corrupted (causal tracing) versus corrupted to clean (resample ablation)—provide insights into the sufficiency or necessity of model components for a given behavior. Clean to corrupted patching identifies activations sufficient for restoring clean performance, even if they are unnecessary due to redundancy, which is particularly informative in OR logic scenarios (Figure 9b, A-OR-B). Conversely, corrupted to clean patching determines the necessary activations for clean performance, which is useful in AND logic scenarios (Figure 9b, A-AND-B).
@@
Activation patching has several limitations, including the effort required to design input templates and counterfactual datasets, the need for human inspection to isolate important subgraphs, and potential second-order effects that can complicate the interpretation of results (Lange et al., 2023) and the hydra effect (McGrath et al., 2023) (see discussion in Section 7.2). Recent advancements aim to address these limitations, such as automated circuit discovery algorithms (Conmy et al., 2023), gradient-based methods for scalable component importance estimation like attribution patching (Nanda, 2023d; Syed et al., 2023), and techniques to mitigate self-repair interferences during analysis (Ferrando & Voita, 2024).

### v3

@@ 4 Core Methods
Mechanistic interpretability (MI) employs various tools, from observational analysis to causal interventions. This section provides a comprehensive overview of these methods, beginning with a taxonomy that categorizes approaches based on their key characteristics (Section 4.1). We then survey observational (Section 4.2), followed by interventional techniques (Section 4.3). Finally, we study their synergistic interplay (Section 4.4). Figure 7 offers a visual summary of the methods and techniques unique to mechanistic interpretability.
Figure 7: Overview of key methods and techniques in mechanistic interpretability research. Observational approaches include structured probes, logit lens variants, and sparse autoencoders (SAEs). Interventional methods, focusing on causal understanding, encompass activation patching variants for uncovering causal mechanisms and causal scrubbing for hypothesis evaluation.
4.1 Taxonomy of Mechanistic Interpretability Methods
We propose a taxonomy based on four key dimensions: causal nature, learning phase, locality, and comprehensiveness (Table 1).
The causal nature of methods ranges from purely observational, which analyze existing representations without direct manipulation, to interventional approaches that actively perturb model components to establish causal relationships. The learning phase dimension distinguishes between post-hoc techniques applied to trained models and intrinsic methods that enhance interpretability during the training process itself.
Locality refers to the scope of analysis, spanning from individual neurons (e.g., feature visualization) to entire model architectures (e.g., causal abstraction). Comprehensiveness varies from partial insights into specific components to holistic explanations of model behavior.
Table 1: Taxonomy of Mechanistic Interpretability Methods
Method Causal Nature Phase Locality Comprehensiveness Key Examples
Feature Visualization Observation Post-hoc Local Partial Zeiler & Fergus (2014)
Zimmermann et al. (2021)
Exemplar methods Observation Post-hoc Local Partial Grosse et al. (2023)
Garde et al. (2023)
Probing Techniques Observation Post-hoc Both Both McGrath et al. (2022)
Gurnee et al. (2023)
Structured Probes Observation Post-hoc Both Both Burns et al. (2023)
Logit Lens Variants Observation Post-hoc Global Partial nostalgebraist (2020)
Belrose et al. (2023)
Sparse Autoencoders Observation Post-hoc Both Comprehensive Cunningham et al. (2024)
Bricken et al. (2023)
Activation Patching Intervention Post-hoc Local Partial Meng et al. (2022a)
Wang et al. (2023)
Path Patching Intervention Post-hoc Both Both Goldowsky-Dill et al. (2023)
Causal Abstraction Intervention Post-hoc Global Comprehensive Geiger et al. (2023a)
Geiger et al. (2023b)
Wu et al. (2023a)
Hypothesis Testing Intervention Post-hoc Global Comprehensive Chan et al. (2022)
Jenner et al. (2023)
Intrinsic Methods Pre/During Global Comprehensive Elhage et al. (2022a)
Liu et al. (2023a)
The categorization is based on the methods’ general tendencies. Some methods can offer local and global or partial and comprehensive interpretability depending on the scope of the analysis and application. Probing techniques can range from local to global and partial to comprehensive; simple linear probes might offer local insights into individual features, while more sophisticated structured probes can uncover global patterns. Sparse autoencoders decompose individual neuron activations (local) but aim to disentangle features across the entire model (global). Path patching extends local interventions to global model understanding by tracing information flow across layers, demonstrating how local perturbations can yield broader insights.
In practice, mechanistic interpretability research involves both method development and their application. When applying methods to understand a model, combining techniques from multiple categories is often necessary and beneficial to build a more comprehensive understanding (Section 4.4).
@@ 4.2 Observation
  1. i.)
    Exemplar methods identify real input examples that highly activate specific neurons or layers. This helps pinpoint influential data points that maximize neuron activation within the neural network (Grosse et al., 2023; Garde et al., 2023; Nanfack et al., 2024).
  2. ii.)
    Feature visualization encompasses techniques that generate synthetic inputs to optimize neuron activation. These visualizations reveal how neurons respond to stimuli and which features are sensitive to (Zeiler & Fergus, 2014; Zimmermann et al., 2021). By inspecting the synthetic inputs that drive neuron behavior, we can hypothesize about the features encoded by those neurons.
@@
Probing (Alain & Bengio, 2016; Hewitt & Manning, 2019) involves training a classifier using the activations of a model, with the classifier’s performance subsequently observed to deduce insights about the model’s behavior and internal representations. However, the probe’s performance may often reflect its own learning capacities more than the actual characteristics of the model’s representations (Belinkov, 2021). This dilemma has led researchers to investigate the ideal balance between the complexity of a probe and its capacity to accurately represent the model’s features (Cao et al., 2021; Voita & Titov, 2020).
The linear representation hypothesis offers a resolution to this issue. Under this hypothesis, the failure of a simple linear probe to detect certain features suggests their absence in the model’s representations. Conversely, suppose a more complex probe succeeds where a simpler one fails. In that case, it implies that the model contains features that a complex function can combine into the target feature but that the target feature itself is not explicitly represented. Thus, the hypothesis implies that using linear probes could suffice in most cases, circumventing the complexity considerations generally associated with probing (Belinkov, 2021).
McGrath et al. (2022) analyzed chess knowledge acquisition in AlphaZero, revealing the emergence of strategic concepts during training. In language models, Gurnee et al. (2023) introduced sparse probing to decode internal neuron activations to understand feature representation and sparsity. They show that early layers use sparse combinations of neurons to represent many features in superposition, while middle layers seem to have dedicated monosemantic neurons for higher-level contextual features.
Probing is limited in drawing causal or behavioral conclusions. Its primarily observational nature focuses on how information is encoded rather than how it is used (see Figure 1), necessitating careful analysis and integration with interventional techniques (Section 4.3), or alternative approaches (Elazar et al., 2021). While in explainable AI, probing has primarily analyzed high-level concepts like linguistic representations (Tenney et al., 2019; Dalvi et al., 2019), MI aims to probe towards uncovering underlying computational processes and functionality. This shift in goals towards uncovering mechanistic computation is a nuanced distinction rather than a clear-cut line between probing in MI and the broader explainability field.
@@
Structured probes aid conceptual interpretability, probing language models for complex features like truth representations. Notably, Burns et al. (2023)’s contrast-consistent search identifies linear projections exhibiting logical consistency in hidden states, contrasting truth values for statements and negations.
However, structured probes face significant challenges in unsupervised probing scenarios. As Farquhar et al. (2023) showed, arbitrary features, not just knowledge-related ones, can satisfy contrast consistency equally well, raising doubts about scalability. For example, the loss may capture simulation of knowledge from hypothesized simulacra within sufficiently powerful language models rather than the models’ true knowledge. Furthermore, Farquhar et al. (2023) demonstrates self-supervised probing methods (like (Burns et al., 2023)) often detect prominent but unintended distractor features in the data. The discovered features are also highly sensitive to prompt choice, and there is no principled way to select prompts that would reliably surface a model’s true knowledge.
@@
Recent work suggests that the essential elements in neural networks are linear combinations of neurons representing features in superposition (Elhage et al., 2022b). To disentangle these features, researchers have developed sparse autoencoders (SAEs), which decompose neural network activations into individual component features (Sharkey et al., 2022b; Cunningham et al., 2024). This process, known as sparse dictionary learning, reconstructs activation vectors as sparse linear combinations of directional vectors within the activation space (Olshausen & Field, 1997).
The theoretical foundations of SAEs are rooted in work on disentangled representations. Whittington et al. (2022) demonstrate that autoencoders can recover ground truth features under conditions of feature sparsity and non-negativity. Furthermore, Garfinkle & Hillar (2019) provides guarantees for the uniqueness and stability of dictionaries for sparse representation, even in the presence of noise. These theoretical underpinnings support SAEs’ ability to uncover true, disentangled features underlying the data distribution.
In practice, SAEs stand out for their simplicity and scalability (Sharkey et al., 2022b). They incorporate sparsity regularization to encourage learning sparse yet meaningful data representations, with the precise tuning of the sparsity penalty on hidden activations critical in dictating the autoencoder’s sparsity level. We provide an overview of the SAE architecture in Figure 8.
SAEs’ dictionary features exhibit higher scores on autointerpretability metrics and increased monosemanticity (Bricken et al., 2023; Cunningham et al., 2024; Sharkey et al., 2022b). They are scalable to state-of-the-art models and can detect safety-relevant features (Templeton et al., 2024), measure feature sparsity (Deng et al., 2023), and interpret reward models in reinforcement learning-based language models (Marks et al., 2023a).
Evaluating SAE quality remains challenging due to the lack of ground-truth interpretable features. Researchers have addressed this through various approaches: Karvonen et al. (2024) proposed using language models trained on chess and Othello transcripts as testbeds, providing natural collections of interpretable features. Sharkey et al. (2022b) constructed a toy model with traceable features, while Makelov et al. (2024); Makelov (2024) compared SAE results with supervised features in large language models to demonstrate their viability.
The versatility of SAEs extends to various neural network architectures. They have been successfully applied to transformer attention layers (Kissane et al., 2024) and convolutional neural networks (Gorton, 2024). Notably, Gorton (2024) applied SAEs to the early vision layers of InceptionV1, uncovering new interpretable features, including additional curve detectors not apparent from examining individual neurons (Cammarata et al., 2020).
In circuit discovery, SAEs have shown particular promise (see also Section 4.4). He et al. (2024) proposed a circuit discovery framework alternative to activation patching (discussed in Section 4.3.1), leveraging dictionary features decomposed from all modules writing to the residual stream. Similarly, O’Neill & Bui (2024) employed discrete sparse autoencoders for discovering interpretable circuits in large language models.
Recent advancements have focused on improving SAE performance and addressing limitations. Rajamanoharan et al. (2024) introduced a gating mechanism to separate the functionalities of determining which directions to use and estimating their magnitudes, mitigating shrinkage – the systematic underestimation of feature activations. An alternative approach by Dunefsky et al. (2024) uses transcoders to faithfully approximate a densely activating MLP layer with a wider, sparsely-activating MLP layer, offering another path to interpretable feature discovery, a type of sparse distillation (slavachalnev, 2024).
Sparse Dictionary Learning Sparse autoencoders (Cunningham et al., 2024) are proposed as a solution to polysemantic neurons. The problem of superposition is mathematically formalized as sparse dictionary learning (Olshausen & Field, 1997) problem to decompose neural network activations into disentangled component features. The goal is to learn a dictionary of vectors $\{{\bm{f}}_{k}\}_{k=1}^{n_{\text{feat}}}\subset{\mathbb{R}}^{d}$ that can represent the unknown, ground truth network features as sparse linear combinations. If successful, the learned dictionary contains monosemantic neurons corresponding to features (Sharkey et al., 2022b). The autoencoder architecture consists of an encoder and a ReLU activation function, expanding the input dimensionality to $d_{\text{hid}}>d_{\text{in}}$. The encoder’s output is given by: ${\bm{h}}$ $=\text{ReLU}({\bm{W}}_{\text{enc}}{\bm{x}}+{\bm{b}}),$ (1) ${\bm{x}}^{\prime}$ $={\bm{W}}_{\text{dec}}{\bm{h}}=\sum_{i=0}^{d_{\text{hid}}-1}h_{i}{\bm{f}}_{i},$ (2) where ${\bm{W}}_{\text{enc}},{\bm{W}}_{\text{dec}}^{\top}\in{\mathbb{R}}^{d_{\text{hid}}\times d_{\text{in}}}$ and ${\bm{b}}\in{\mathbb{R}}^{d_{\text{hid}}}$. The parameter matrix ${\bm{W}}_{\text{dec}}$ forms the feature dictionary, with rows ${\bm{f}}_{i}$ as dictionary features. The autoencoder is trained to minimize the loss, where the $L^{1}$ penalty on $\bm{h}$} encourages sparse reconstructions using the dictionary features, $\mathcal{L}({\bm{x}})=|{\bm{x}}-{\bm{x}}^{\prime}|_{2}^{2}+\alpha|{\bm{h}}|_{1}.$ (3) [Uncaptioned image] Figure 8: Illustration of a sparse autoencoder applied to the MLP layer activations, consisting of an encoder that increases dimensionality while emphasizing sparse representations and a decoder that reconstructs the original activations using the learned feature dictionary.
@@ 4.3 Intervention
The theory of causality (Pearl, 2009) provides a mathematically precise framework for mechanistic interpretability, offering a rigorous approach to understanding high-level semantics in neural representations (Geiger et al., 2023a). By treating neural networks as causal models, with their compute graphs serving as causal graphs, researchers can perform precise interventions and examine the roles of individual parameters (Mueller et al., 2024). This causal perspective on interpretability has led to the development of various intervention techniques, including activation patching (Section 4.3.1), causal abstraction (Section 4.3.2), and hypothesis testing methods (Section 4.3.3).
@@ 4.3.1 Activation Patching
Figure 9: (a) Activation patching in a transformer model. Left: The model processes the clean input "Colosseum in Rome," caching the latent activations (step i.). Right: The model runs with the corrupted input "Eiffel Tower in Paris" (step ii.). The pink arrow shows an MLP layer activation (green diamond) patched from the clean run into the corrupted run (step iii.). This causes the prediction to change from "Paris" to "Rome," demonstrating how the significance of the patched component is determined (step iv.). By comparing these carefully selected inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the location prediction behavior. (b) Activation patching directions: Top: Patching corrupted activations (orange) into clean circuits (turquoise) reveals sufficient components for identifying OR logic scenarios. Bottom: Patching clean activations (green) into corrupted circuits (orange) reveals necessary components that are useful for identifying AND logic scenarios. The AND and OR gates demonstrate how these patching directions uncover different logical relationships between model components.
@@
The standard protocol for activation patching (Figure 9a) involves:
  1. step i.
    Running the model with a clean input and caching the latent activations;
  2. step ii.
    Executing the model with a corrupted input;
  3. step iii.
    Re-running the model with the corrupted input but substituting specific activations with those from the clean cache; and
  4. step iv.
    Determining significance by observing the variations in the model’s output during the third step, thereby highlighting the importance of the replaced components.
This process relies on comparing pairs of inputs: a clean input, which triggers the desired behavior, and a corrupted input, which is identical to the clean one except for critical differences that prevent the behavior. By carefully selecting these inputs, researchers can control for confounding circuitry and isolate the specific circuit responsible for the behavior.
Differences in patching direction – clean to corrupted (causal tracing) versus corrupted to clean (resample ablation) – provide insights into the sufficiency or necessity of model components for a given behavior. Clean to corrupted patching identifies activations sufficient for restoring clean performance, even if they are unnecessary due to redundancy, which is particularly informative in OR logic scenarios (Figure 9b, OR gate). Conversely, corrupted to clean patching determines the necessary activations for clean performance, which is useful in AND logic scenarios (Figure 9b, AND gate).
@@
Recently, Ghandeharioun et al. (2024) introduced patchscopes, a framework that unifies and extends activation patching techniques: using the model’s text generation to explain internal representations, it enables more flexible interventions across various interpretability tasks, improving early layer inspection and allowing for cross-model analysis.
@@
Activation patching has several limitations, including the effort required to design input templates and counterfactual datasets, the need for human inspection to isolate important subgraphs, and potential second-order effects that can complicate the interpretation of results (Lange et al., 2023) and the hydra effect (McGrath et al., 2023; Rushing & Nanda, 2024) (see discussion in Section 7.2). Recent advancements aim to address these limitations, such as automated circuit discovery algorithms (Conmy et al., 2023), gradient-based methods for scalable component importance estimation like attribution patching (Nanda, 2023d; Syed et al., 2023), and techniques to mitigate self-repair interference during analysis (Ferrando & Voita, 2024).